OpenStudy (anonymous):

Write an equation in y = mx + b form for the table. Gallons of gas used 2 4 6 8 Miles driven 52 104 156 208

OpenStudy (anonymous):

m = (104-52)/2 = 26 52 = 2*26 + b -> b=0 y=26*x

OpenStudy (anonymous):

y=26x